# CamSAM2: Segment Anything Accurately in Camouflaged Videos

### v2

@@ 3.6 Training and Inference Strategies
Table 2: Detailed comparisons between CamSAM2 and SAM2 on MoCA-Mask. CamSAM2 consistently outperforms SAM2 for all considered prompt types and backbones. Improvements of CamSAM2 over SAM2 are shown in dark green.
EGNet [60] ResNet-50 111.7 - 61.9 29.8 4.4 35.0 66.6 32.4 24.3
BASNet [37] ResNet-50 87.1 - 63.9 34.9 5.4 39.4 77.3 39.3 29.3
CPD [48] ResNet-50 47.9 - 62.2 28.9 4.9 35.7 66.7 33.0 23.9
PraNet [11] ResNet-50 32.6 - 62.9 35.2 4.2 39.7 76.3 37.8 29.0
SINet [10] ResNet-50 48.9 - 63.6 34.6 4.1 39.5 77.5 38.1 28.3
SINet-V2 [12] Res2Net-50 27.0 - 65.3 38.2 3.9 43.2 76.2 41.3 31.8
PNS-Net [16] ResNet-50 142.9 - 65.5 32.5 4.8 41.7 67.3 38.4 29.0
RCRNet [52] ResNet-50 53.8 - 62.7 28.7 4.8 32.8 66.6 30.9 22.9
MG [53] VGG 4.8 - 59.4 33.6 5.9 37.5 69.2 36.8 26.8
SLT-Net-LT [7] PVTv2-B5 82.3 - 69.6 48.1 3.0 52.4 84.5 49.3 40.2
ZoomNeXt [33] PVTv2-B5 84.8 - 75.7 59.3 2.0 63.1 86.5 59.9 51.0
SAM2 [39] Hiera-T 38.9 1-click 75.7 58.3 3.3 62.2 81.4 59.2 48.9
CamSAM2 Hiera-T 39.4 1-click 77.1 62.2 3.2 68.1 83.9 62.6 50.7
SAM2 [39] Hiera-T 38.9 box 85.4 77.3 1.7 79.5 95.1 77.8 66.7
CamSAM2 Hiera-T 39.4 box 87.2 79.5 1.3 81.4 96.3 79.6 69.2
Table 3: Comparisons between our method and existing approaches on CAD. Our approach clearly outperforms existing methods.
SUN-SEG-Easy
SAM2 [39] 83.4 71.6 83.0 73.6
CamSAM2 88.3 82.6 93.4 84.3
SUN-SEG-Hard
SAM2 [39] 75.5 58.4 73.4 61.0
CamSAM2 86.4 78.2 91.2 80.6
4 Experiments
4.1 Experimental Setup
Datasets.  Our experiments are conducted on three video datasets: two popular camouflaged animal datasets, MoCA-Mask [7] and CAD [3], and one camouflaged medical dataset, SUN-SEG [17]. The pioneering Moving Camouflaged Animals dataset (MoCA) [21] comprises 37K frames from 141 YouTube video sequences. The dataset MoCA-Mask is reorganized from the MoCA, containing 71 video sequences with 19,313 frames for training and 16 video sequences with 3,626 frames for testing, respectively, with pixel-wise ground-truth masks on every five frames. It also generates a MoCA-Mask-Pseudo dataset, which contains pseudo masks for unlabeled frames with a bidirectional optical-flow-based consistency check strategy. The Camouflaged Animal Dataset (CAD) includes 9 short videos in total that have 181 hand-labeled masks on every five frames. SUN-SEG is the largest benchmark for video polyp segmentation, derived from SUN-database [30]. It consists of a training set with 112 clips (19,544 frames) and two test sets: SUN-SEG-Easy, containing 119 clips (17,070 frames), and SUN-SEG-Hard, comprising 54 clips (12,522 frames).
Implementation Details.  The proposed CamSAM2 is implemented with PyTorch [35]. CamSAM2 is initialized with the parameters of SAM2. We freeze all parameters used in SAM2 and initialize other parameters randomly. We set $betas=(0.9,~{}0.999)$ for the optimizer Adam and use the initial learning rate of 1e-3 with the StepLR of 10. We train CamSAM2 on 4 NVIDIA RTX 4090 GPUs for 15 epochs. For camouflaged animal segmentation, we train the model using the MoCA-Mask-Pseudo training set and evaluate it on the MoCA-Mask test set and CAD. During inference for SAM2 and CamSAM2, we apply the 1-click, box, and mask prompts only on the first frame of each video. For camouflaged polyp segmentation, we train the model using the SUN-SEG training set and perform inference using mask prompt on the first frame of each video on the SUN-SEG-Easy and SUN-SEG-Hard test sets.
Evaluation Metrics.  We adopt seven evaluation metrics to measure the quality of predicted pixel-wise masks: S-measure ($S_{m}$[8], F-measure ($F_{\beta}$[1], weighted F-measure ($F_{\beta}^{\omega}$[28], mean absolute error (MAE) [36], E-measure ($E_{m}$[9], mean Dice (mDice), and mean IoU (mIoU).
4.2 Experimental Results
Figure 4: Visualization of the attention maps obtained from SAM2 and CamSAM2 using 1-click point prompt with the Hiera-T backbone. The attention map is extracted from the last token-to-image cross-attention in the mask decoder. From top to bottom: the input frames, attention map of SAM2’s mask token and the image embedding, attention map of the decamouflaged token and the image embedding, and ground-truth masks. The higher attention regions are indicated by warmer colors.
Results on MoCA-Mask.Tab. 1 compares the performance of three promptable methods. CamSAM2 clearly outperforms SAM-PM and SAM2; even with the 1-click prompt, CamSAM2 still outperforms SAM-PM, which uses mask prompt on the first frame. The promptable methods clearly outperform other models on the MoCA-Mask, highlighting the strength of integrating prompt-based strategies for VCOS and demonstrating the potential of promptable methods to excel in scenarios where camouflaged objects are particularly challenging to segment and track.
Tab. 2 presents a detailed comparison between SAM2 and CamSAM2 on the MoCA-Mask dataset, across different prompt types (1-click, box, and mask) with Hiera-T and Hiera-S backbones. CamSAM2 demonstrates consistent improvements over SAM2 across all prompt types and backbones. With a 1-click prompt, it achieves mDice/mIoU gains of 12.2/9.8 (Hiera-T) and 13.0/12.0 (Hiera-S), demonstrating its effectiveness in segmenting camouflaged objects with minimal input. For a box prompt, CamSAM2 improves mIoU by 2.5 (Hiera-T) and 1.6 (Hiera-S). With a mask prompt, it achieves 2.6 and 0.9 mIoU gains for Hiera-T and Hiera-S, respectively.
Fig. 3 shows qualitative results for two video clips of MoCA-Mask from SAM2 and CamSAM2 using 1-click prompts with the Hiera-T as the backbone. Fig. 4 presents a comparative visualization of the attention maps generated by SAM2 and CamSAM2 for various objects. The attention maps are extracted from the last token-to-image cross attention layer in the mask decoder. The token serves as the query, and the image embedding serves as the key and value. CamSAM2 demonstrates superior attention precision over SAM2, with a larger activated region and stronger activations around target objects, validating the effectiveness of our proposed methods in enhancing VCOS quality.
Notably, CamSAM2 introduces only a marginal increase in parameters (0.5M). Despite this minimal increase, CamSAM2 achieves significant improvements while keeping all of SAM2’s parameters unchanged, fully inheriting SAM2’s capability for segmenting and tracking common objects in natural scenes.
Results on CAD.  We evaluate the zero-shot performance of CamSAM2 and SAM2 on the CAD dataset using Hiera-T and Hiera-S backbones with point and box prompts, as shown in Tab. 3. CamSAM2 demonstrates notable improvements over SAM2, particularly in the 1-click prompt setting, where it gains 3.4 and 1.8 in mDice and mIoU, highlighting CamSAM2’s enhanced capability when only minimal guidance is available. In the box prompt setting, CamSAM2 also shows clear gains, with an increase of 1.8 and 2.5 in mDice and mIoU. These observations indicate that CamSAM2 outperforms SAM2 in zero-shot scenarios, underscoring its effectiveness and suitability for practical segmentation tasks that require minimal user input.
Results on SUN-SEG.Tab. 4 presents the performance comparison on the SUN-SEG dataset, showing that CamSAM2 consistently outperforms SAM2 across all metrics on both SUN-SEG-Easy and SUN-SEG-Hard test sets. Notably, CamSAM2 achieves substantial improvements. Specifically, mDice improves by 10.7 on SUN-SEG-Easy, rising from 73.6 to 84.3, and by 19.6 on SUN-SEG-Hard, increasing from 61.0 to 80.6, demonstrating its effectiveness in segmenting camouflaged polyps. Based on the results above, our method significantly enhances SAM2 across different camouflaged scenarios, demonstrating its effectiveness and broad applicability in VCOS.
 Decam.
Token
 IOF  EOF  OPG  mDice $\uparrow$  mIoU $\uparrow$
52.1 44.8
54.9 47.0
55.2 47.5
55.9 47.9
64.3 54.6
4.3 Ablation Studies
To understand the impact of individual components in CamSAM2, we conducted ablation studies on the MoCA-Mask test set using the Hiera-T backbone and a 1-click point prompt. The goal is to measure the contributions of key components, including the decamouflaged token, IOF, EOF, and OPG, on segmentation performance. Additionally, we evaluated the effects of different distance metrics and prototype numbers in the OPG process. We analyze the results below.
[…]

### v3

@@ A.3 Comparative analysis across point prompt with different numbers of clicks
SAM2 5-click 65.9 56.7
CamSAM2 68.1
(+2.2)
58.6
(+1.9)
Hiera-S
SAM2 1-click 54.9 46.7
CamSAM2 68.0
(+13.1)
58.8
(+12.1)
SAM2 2-click 56.8 47.8
CamSAM2 63.6
(+6.8)
53.7
(+5.9)
SAM2 3-click 71.7 61.7
CamSAM2 73.6
(+1.9)
62.9
(+1.2)
SAM2 5-click 68.5 58.4
CamSAM2 73.9
(+5.4)
63.0
(+4.6)
A.4 Comparison with SAM2-Adapter
Table 10: Comparison between SAM2-Adapter and CamSAM2 with Hiera-L as the backbone.
Model Prompt $S_{m}\uparrow$ $F_{\beta}^{\omega}\uparrow$ MAE $\downarrow$ $F_{\beta}\uparrow$ $E_{m}\uparrow$ mDice $\uparrow$ mIoU $\uparrow$
SAM2-Adapter [chen2024sam2adapterevaluatingadapting] - 68.4 38.7 0.9 43.0 80.0 43.2 35.8
CamSAM2 1-click 82.2 71.8 0.6 73.4 92.1 73.5 63.3
To provide a fair and comprehensive comparison, we evaluate CamSAM2 alongside SAM2-Adapter [chen2024sam2adapterevaluatingadapting]. SAM2-Adapter is an approach to adapt SAM2 to downstream tasks and achieve enhanced performance. This method effectively integrates task-specific knowledge with the general knowledge learned by the model. During this process, only the adaptor layers and the mask decoder are trained, while the rest of the image encoder is kept frozen. Since the official SAM2-Adapter implementation uses Hiera-L as the backbone, we directly compare on the same backbone to avoid unintended modifications to their codebase. We re-train the SAM2-Adapter on the MoCA-Mask dataset using their provided settings. Results are reported in Tab.˜10. Importantly, CamSAM2 and SAM2-Adapter represent two fundamentally different design philosophies for extending SAM2:
Evaluation granularity.
SAM2-Adapter operates at the image level, making per-frame predictions independently. In contrast, CamSAM2 is designed for video level inference, where only the first frame receives a prompt, and subsequent frames are segmented using memory-augmented features that evolve over time. This enables prototype refinement and accumulated improvements across frames, which is a capability SAM2-Adapter lacks.
Prompt usage.
SAM2-Adapter relies on a learned prompt injected into different stages in the image encoder at every image during inference. In contrast, CamSAM2 follows the original SAM2 design as a promptable segmentation model, accepting a single external prompt (e.g., a click) on the first frame and leveraging internal mechanisms to propagate the segmentation over time. While our usage of ground-truth-derived prompts may raise fairness concerns, this is a common and practical setting in semi-supervised video segmentation and is consistent with SAM2’s prompt-driven paradigm.
In summary, CamSAM2 and SAM2-Adapter offer two distinct solutions. CamSAM2 enhances SAM2 for video by leveraging memory and minimal user guidance, while SAM2-Adapter adapts SAM2 to images through internal prompt learning. Although they address different use cases, our results demonstrate that CamSAM2 provides a more effective and temporally coherent framework for VCOS, maintaining compatibility with SAM2’s promptable design without modifying its parameters.
A.5 Additional training and inferencing details
As we mentioned in the §4.1, we adopt a combined loss of binary cross-entropy (BCE) and dice loss for mask predictions across the entire video sequence. This loss applies to both the original SAM2 mask logits $\mathbf{R}_{i}$ and our CamSAM2 mask logits $\mathbf{R}_{i}^{c}$. Although the original SAM2 parameters remain frozen during training, we still apply supervision to its outputs. The reason is that the introduction of our learnable decamouflaged token changes the self-attention dynamics in the mask decoder: this token is concatenated with the original output tokens and jointly participates in self-attention and cross-attention layers (see Fig.˜2(a) in the main paper). Consequently, the outputs $\mathbf{R}_{i}$ are no longer guaranteed to be identical to those of the original SAM2. By supervising them, we ensure that SAM2’s predictions remain aligned with the segmentation intent and are not perturbed by the token interactions.
Inspired by HQ-SAM [ke2023segment], we adopt a simple post-processing strategy where the final mask prediction is obtained by averaging SAM2’s mask logits $R_{t}$ and the CamSAM2 mask logits $R_{t}^{c}$, which is calculated by the SAM2 mask token and the decamouflaged token. This strategy integrates SAM2’s global representation with the fine-grained, camouflage-aware features extracted by our proposed modules. Despite its simplicity, this approach effectively unifies global generalization and fine-grained object sensitivity.
Tab.˜11 reports mIoU scores on the Hiera-T backbone. The first column shows the original SAM2 performance. The second and third columns report the mask predictions from the SAM2 token and the decamouflaged token within CamSAM2, respectively. The final column presents the result of averaging both. Each individual output mask already outperforms the SAM2 baseline, and their combination leads to further improvements across all prompt types, validating the effectiveness of our averaging strategy.
Table 11: Comparison of SAM2, individual token predictions within CamSAM2, and their combination. Results are reported as mIoU on the Hiera-T backbone across three prompt types. “SAM2 baseline” refers to the vanilla SAM2 prediction. “SAM2 token” denotes the prediction generated from the SAM2 mask token, while “decamouflaged token” refers to the prediction from the decamouflaged token after integrating features from our proposed IOF, EOF, and OPG modules. “CamSAM2” is the final result obtained by averaging the mask logits calculated by the SAM2 token and the decamouflaged token.
Prompt SAM2 baseline SAM2 token decamouflaged token CamSAM2
1-click 44.8 53.3 53.1 54.6
box 62.3 63.9 64.1 64.8
mask 67.9 70.3 68.4 70.5
Notably, the mask of the SAM2 output token within CamSAM2 yields better results than in the original SAM2 model, despite having identical parameters and no additional fine-tuning. This improvement stems from two key factors. First, all output tokens, including the SAM2 mask tokens and the decamouflaged token, are jointly processed in the same transformer, where they interact through a self-attention layer and different cross-attention layers. Second, SAM2’s architecture conditions the current frame’s feature $F_{t}^{mem}$ on the previously predicted mask stored in memory, enriching the encoder’s output over time. This mutual influence further enhances their individual representational capacity.
A.6 First frame performance analysis
We evaluate CamSAM2 on the first frame of each video clip on the MoCA-Mask dataset using a 1-click point prompt with Hiera-T as the backbone. This setting isolates the model’s ability to handle camouflaged object segmentation, as on the first frame, the OPG module is naturally inactive, since there are no prior frames from which to extract prototypes. This makes the first frame an ideal testbed for evaluating the model’s segmentation capability only without temporal information.
Table 12: First frame comparison between SAM2 and CamSAM2 with the point prompt.
Model $S_{m}\uparrow$ $F_{\beta}^{\omega}\uparrow$ MAE $\downarrow$ $F_{\beta}\uparrow$ $E_{m}\uparrow$ mDice $\uparrow$ mIoU $\uparrow$
SAM2 72.2 56.7 7.4 58.1 75.9 58.6 51.1
CamSAM2 73.0 59.4 7.2 60.8 80.7 61.4 52.6
As shown in Tab.˜12, CamSAM2 consistently outperforms SAM2 across all metrics on the first frame. This improvement is particularly meaningful, as it demonstrates that CamSAM2 enhances SAM2’s segmentation capability even in the absence of temporal cues, laying a stronger foundation for subsequent predictions. Since CamSAM2 stores the predicted mask of each frame into memory and uses it to generate object prototypes, these prototypes can progressively enhance segmentation in future frames within EOF, resulting in cumulative performance gains over time, which shows that CamSAM2 further enhances the tracking ability of SAM2 for the VCOS task.
A.7 Efficiency analysis of OPG
Table 13: Average runtime per frame and overhead breakdown of CamSAM2.
Component Avg Time (ms) Relative Percentage
SAM2 83.7 100.0%
FPS 3.5 4.2%
k-means 1.0 1.2%
OPG 4.6 5.5%
CamSAM2 89.7 107.2%
To assess the runtime efficiency of CamSAM2, we profile its per-frame latency using the Hiera-T backbone on a representative 106-frame camouflaged video from the MoCA-Mask test set. Tab. 13 reports the average runtime breakdown and compares CamSAM2 against the baseline SAM2. Overall, CamSAM2 introduces only 6 ms (7.2%) additional latency per frame relative to SAM2, demonstrating that the method remains efficient. The OPG module, which consists of FPS sampling and 1-iter $k$-means clustering, accounts for the majority of the runtime overhead. All other added components, including IOF, EOF, and associated processing, together contribute approximately 1.4 ms (1.7%) per frame.
We profile the computational complexity in Tab. 14. CamSAM2 introduces an additional 14.4 GFLOPs, which is independent of different backbone architectures. Based on the Hiera-T backbone, the total FLOPs increase from 137.2 GFLOPs to 151.5 GFLOPs, which corresponds to approximately a 10% increase. This added cost results in a significant performance gain of 9.8 mIoU with point prompts. Moreover, when using the Hiera-L backbone, which image encoder requires 810 GFLOPs, the same 14.4 GFLOPs overhead accounts for only a 1.7% increase in total FLOPs. Even with such a small relative increase, CamSAM2 still achieves a notable 7.5 mIoU gain.
Table 14: FLOPs comparison between SAM2 and CamSAM2.
Module FLOPs (G)
Image encoder 103.0
Memory encoder 5.0
Memory attention 27.4
Mask decoder 1.8
SAM2 137.2
CamSAM2 extra 14.4
CamSAM2 151.5
These findings confirm that CamSAM2 remains efficient and is suitable for practical deployment, even in time-sensitive VCOS applications.
A.8 Additional analysis and ablation studies
Number of Decamouflaged Tokens.
We conduct an ablation study on the number of decamouflaged tokens to evaluate its effect on representation quality. As shown in Tab. 15, increasing the number of tokens from 1 to 3 results in a comparable performance in mIoU. This indicates that adding more tokens does not structurally improve the representation quality, but may introduce redundancy and additional computational cost.
Table 15: Impact of using different numbers of decamouflaged tokens.
Model Number of Decam. Token mIoU
SAM2 - 44.8
CamSAM2 3 53.3
CamSAM2 1 54.6
Sampling and Clustering Strategies.
To evaluate the robustness and generality of the OPG module, we experiment with different strategies for sampling and clustering. Specifically, we compare (i) Farthest Point Sampling (FPS) versus Average Sampling for selecting initial cluster centers, and (ii) $k$-means versus Gaussian Mixture Model (GMM) for clustering. Results are reported in Tab. 16. All combinations yield reasonable performance, confirming the flexibility of OPG. Notably, FPS better preserves spatial diversity among selected features, leading to more representative prototypes.
Table 16: Impact of different sampling and clustering strategies in OPG.
Sampling Clustering mIoU
Average Sampling $k$-means 52.1
Average Sampling GMM 52.7
FPS GMM 53.6
FPS $k$-means 54.6
These findings show that OPG is not overly sensitive to specific implementation choices, while FPS sampling in combination with $k$-means clustering offers the best trade-off between representativeness and segmentation accuracy.
A.9 Architecture toggle between CamSAM2 and SAM2
Figure 5: Illustration of the architecture toggle. The toggle switch enables or disables the proposed modules for VCOS containing the decamouflaged token, IOF, EOF, and OPG. Modules and flows in dashed lines indicate the disabled state.
To better illustrate the relationship between CamSAM2 and the original SAM2 pipeline, Fig.˜5 presents a side-by-side comparison. When the toggle is set to OFF, the proposed modules are completely bypassed, and the pipeline reverts to the SAM2 architecture. Importantly, since all SAM2 parameters are fixed during the training of CamSAM2, the model behaves identically to SAM2 in this mode, preserving its original structure and performance across various tasks. This design ensures compatibility and flexibility, allowing seamless integration without disrupting SAM2’s performance.
A.10 Limitation
CamSAM2 is built upon the SAM2 architecture, inheriting both its strengths and limitations. While our proposed modifications significantly improve SAM2’s performance in VCOS tasks, we do not alter the core architecture of SAM2. As a result, CamSAM2 retains several of SAM2’s known limitations. Specifically, it does not address SAM2’s challenges in handling shot changes and long occlusions. Future work may consider integrating explicit motion information or relational modeling to further improve the robustness of the model.
A.11 Interpreting prototypes via cosine similarity maps
To improve interpretability, we visualize the cosine similarity between the object prototype from preceding frames and the current-frame feature map of two selected videos. As shown in Fig. 6, the similarity maps consistently highlight semantically meaningful object regions while suppressing the background, indicating that the prototypes abstract object-aware semantics and remain temporally consistent rather than drifting to spurious parts.
Figure 6: The cosine similarity map between the preceding frames prototype and the current frame feature map.
A.12 More qualitative result visualization
Figure 7: Qualitative comparisons between SAM2 and CamSAM2 using mask prompt with the Hiera-T backbone on two video clips of SUN-SEG-Hard. From top to bottom: the input frames, SAM2’s results, CamSAM2’s results, and ground-truth masks. CamSAM2 demonstrates improved accuracy in segmenting camouflaged polyps, as shown by the circles. Best viewed in color.
Fig.˜7 presents two qualitative example video clips of SUN-SEG-Hard, comparing the segmentation performance of SAM2 and CamSAM2 using mask prompts with the Hiera-T backbone. Compared with SAM2, CamSAM2 can segment polyps more accurately, including more precise regions and boundaries.
Fig.˜8 and Fig.˜9 present more comprehensive qualitative results for six video clips in MoCA-Mask, comparing the segmentation performance of SAM2 and CamSAM2 using 1-click point prompts with the Hiera-T backbone. The results clearly demonstrate CamSAM2’s advantages in segmenting camouflaged objects in videos. For instance, in the first (stick insect) and second (rusty spotted cat) examples from Fig.˜8, SAM2 is able to segment parts of the camouflaged objects; however, CamSAM2 significantly outperforms by successfully segmenting larger and more complete regions of the objects. Furthermore, in the third (pygmy seahorse) example, SAM2 produces an incorrect segmentation, failing to segment the camouflaged object accurately, while CamSAM2 still manages to segment the correct regions of the object. These findings highlight CamSAM2’s ability to both enhance accuracy in cases where SAM2 performs moderately well and to maintain reliable segmentation performance in challenging scenarios where SAM2 fails. This underscores CamSAM2’s robustness and effectiveness in tackling diverse and complex tasks in VCOS.
Figure 8: More qualitative comparisons between SAM2 and CamSAM2 using 1-click point prompt with the Hiera-T backbone on three video clips. From top to bottom: the input frames, SAM2’s results, CamSAM2’s results, and ground-truth masks. CamSAM2 demonstrates improved accuracy in segmenting camouflaged objects, as shown by the circles. Best viewed in color.
Figure 9: More qualitative comparisons between SAM2 and CamSAM2 using 1-click point prompt with the Hiera-T backbone on three video clips. Best viewed in color.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction provide a comprehensive overview of the background and motivation of this study, accurately reflecting its main contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discussed the limitation in the supplementary material.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: No theoretical result is included in this paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All information regarding the key contribution of this paper including architectural, data, and experimental configurations, has been fully disclosed.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The datasets used in this paper are publicly available. The code is released at github.com/zhoustan/CamSAM2.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All details have been discussed in §4.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We only provide baseline results for future reference. We will release code to reproduce all our reported results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We discussed the compute resources in §4.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in our paper complies in all respects with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discussed the potential positive impacts in §5.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
[…]